A tugboat pulls a ship with a constant net horizontal force of 5.00 × 103 N
and causes the ship to move through a harbor. How much work is done on
the ship if it moves a distance of 3.00 km?
1
Expert's answer
2020-04-09T09:30:50-0400

As per the given question,

Net horizontal force"= 5.00 \u00d7 10^3 N"

Distance moved by the ship (d)"=3km=3000m"

Hence, net work done "(W)= F.d"

"\Rightarrow W= 5.0\times 10^3\times 3\times 10^3J"

"\Rightarrow W=15\times 10^6J"